I have a question on how to do this problem, |x/2|=3.
I don't kmow how to do it the absolute value way.
I mean would it be negative x and 2 or would it be just negative x or 2.

Answer by 2897696(96)

remember, when an absolute has and equal sign, first solve the absolute values positive: x/2=3 ==> x=6 (meaning leave answer(3) positive)
then solve it negative: x/2=-3 ==> x=-6 (meaning make answer(3) negative)
so the answer to the absolute value is 6 and -6
ANSWER: 6 and -6
try these questions:
1)|x-3| =4
2)|2x+7|=9
3)|v-10|=2